OpenStudy (anonymous):

Write the equation of a line parallel to the line 2x + 3y = 5 and passes through the point (9, -3)

OpenStudy (anonymous):

if two lines are parellel to each other, they have the same slope m

OpenStudy (anonymous):

try to find the slope of the given equation and use y=mx+b and plug in (9.-3) to get the value of b. problem solved ;)

OpenStudy (texaschic101):

do you know how to find the slope ?

OpenStudy (texaschic101):

y = mx + b -- the slope is m so put your equation in y = mx + b form and m is your slope. Give it a try